some ethylene glycol is added to your car's cooling system along with 5.0 kg. of water. If the freezing point of the water glycol solution is -15.0 degree celsius what mass of HOCH2CH2OH must have been added?


1
Expert's answer
2021-11-22T17:08:03-0500

Taking Kf of water as "=1.86K\>kgmol^{-1}"


"-\Delta\>T_f=K_f×M"


"0-(-15)=1.86\u00d7M"


"M=\frac{15}{1.86}=8.0645\>moles"


Molar mass of ethylene glycol

= 62g

Mass added in 5kg of water

=8.0645×62×5

=2499.995g

=2.5kg